OpenStudy (anonymous):

At Pizitz, 48 students take math team, 20 are in Scholars Bowl and 78 take band. 6 students are in both band and Scholars Bowl, 8 take both Scholars Bowl and math team, 21 take both math team and band, and 3 take all three activities. What percent of math team students only participate in math team? please show me the steps.

OpenStudy (anonymous):

1. add all of the students that take math team with something else 8+21+3=32 2. subtract the students that take math team with something else from the students who take math team 48-32=16 3. take the kids who take just math team over all of the math team students and multiply it by 100. (16/48)100= 33 and 1/3 %

OpenStudy (anonymous):

Thank you so much !